# Adapting Large Language Models for Document-Level Machine Translation

### 12 Jan 2024 v1

@@ 5 Main Results
In our results presented in Table 2, we observe that gpt-4-turbo and gpt-3.5-turbo significantly outshine all other models in performance. Notably, the NLLB variants, which are trained on vast amount of parallel sentence pairs, also demonstrate superior performance among specialized machine translation (MT) models. In the context of DocMT, Doc2Doc-mT5 models still outperform our LLM-based DocMT models for translations from English to other languages when evaluated using standard MT metrics. Conversely, for translations from other languages to English, our LLM-based DocMT models outperform Doc2Doc-mT5 in $\mu_{s\textsc{BLEU}}$ and $\mu_{d\textsc{BLEU}}$ metrics, while the Doc2Doc-mT5 models maintain a superior performance in $\mu_{\textsc{COMET}}$.
@@
As indicated in Table 2, our models incorporating LoRA typically outperform fully fine-tuned (FFT) LLMs. However, an exception is observed where V-7B-fft outperforms V-7B-LoRA in translating from other languages to English. This discrepancy is likely attributable to overfitting. In scenarios of extensive fine-tuning with a large corpus of parallel documents, the full fine-tuning of all parameters often leads to rapid overfitting on the training dataset. In contrast, the parameter-efficient fine-tuning approach, exemplified by LoRA, updates only a select number of parameters, effectively preventing the models from overfitting the training set. Furthermore, we observe that the L-7B and V-7B models exhibit comparable performance, suggesting that initializing the models with instruction-tuned models does not necessarily enhance task-specific performance.
@@
We present the results for the translation tasks from other languages to English in Figure 2. Regarding the readability of the figures, we present only the results provided by our models using LoRA. Our LLM-based DocMT models exhibit superior performance, sometimes even surpassing gpt-4-turbo in certain translation tasks. However, they fail completely in others. A manual review of translation tasks where our LLM-based DocMT models fail reveals that the primary cause of failure is off-target translation. We provide an in-depth analysis of the off-target translation problem in Section 6.1. A complete breakdown of the results is in Appendix D.
002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8002,00020002{,}0002 , 0002,20022002{,}2002 , 2002,40024002{,}4002 , 400MistranslationOvertranslationUndertranslationAdditionOmissionGrammarUnclear referenceCohesionCoherenceInconsistent styleMultiple terms in translation1,39513951{,}3951 , 3957177177177171,22412241{,}2241 , 224840840840840484484484484234234234234646464647097097097094544544544547257257257253463463463461,35613561{,}3561 , 3567177177177171,35813581{,}3581 , 3588638638638636166166166162662662662661021021021027277277277275655655655657717717717713393393393391,71217121{,}7121 , 712715715715715977977977977245245245245417417417417486486486486808080808048048048045925925925928248248248242732732732732,00220022{,}0022 , 0028368368368361,55115511{,}5511 , 551586586586586868868868868474474474474206206206206933933933933717717717717854854854854386386386386# of occurences
@@ 6 Analyses
In this section, we investigate the off-target problem (Section 6.1), leverage gpt-4-turbo to analyze the translation errors (Section 6.2), explore the scaling law of parallel documents (Section 6.3), conduct additional evaluations on recent test sets from WMT2023 (Section 6.4), and examine crosslingual transfer (Section 6.5).
@@
6.1 Off-Target Translation
Figure 2 demonstrates that our LLM-based DocMT models excel in certain translation tasks but fail in others. The key reason for these failures is the off-target translation issue, where translations are in the incorrect language. In this section, we utilize the fasttext library (Bojanowski et al., 2017) to determine the language of the translations produced by our models. [^6] https://huggingface.co/facebook/fasttext-language-identification We then calculate the off-target rate, which is the proportion of translations that are off-target, to quantify this problem.
We display the off-target rates of our models for translations from various languages to English in Table 3. In tasks where our models fail entirely, the off-target rate can reach up to $98.3\%$. Remarkably, only B-7B-LoRA maintains a low off-target rate across all language pairs. These results highlight the primary reason for the translation failures in LLM-based DocMT models and provide valuable insights for future research. The complete off-target rates are detailed in Appendix E.
6.2 Translation Errors
@@
To comprehensively understand the translation capabilities of our LLM-based DocMT models, we select specific error types from the Multidimensional Quality Metrics (MQM) framework (Burchardt, 2013). Kocmi and Federmann (2023) demonstrate gpt-4 is capable of identifying error spans and achieving state-of-the-art MT evaluation accuracy, so we leverage gpt-4-turbo to analyze the translation errors of the text translated by these models. In this study, due to resource limitations, we focus on translations performed by four models: L-7B-LoRA, L-7B-fft, Doc2Doc-mT5-1.2B, and GoogleTrans. These models are evaluated on translation tasks from English to German, Romanian, and Chinese. The prompt used to identify errors is detailed in Appendix C.
In Figure 3, we display the frequency of various error types observed. Notably, most errors are confined to individual sentences. Since we instructed gpt-4-turbo to focus on the relationship between the current sentence and its context, a significant number of context-dependent errors are identified. Despite similar scores in metrics like $s$BLEU, $d$BLEU, and COMET among the evaluated models, our LLM-based DocMT models (L-7B-LoRA and L-7B-fft) exhibit fewer context-independent and context-dependent errors. This finding highlights a limitation in current evaluation metrics, suggesting they may not adequately assess translations at the document level. It also indicates that fine-tuning LLMs for machine translation is a promising research direction, particularly for improving document-level translation quality.
1102030405060708090100404040406060606080808080100100100100Percentage (%) of training dataCOMET
@@
6.3 Scaling Law of Parallel Documents
In this section, we delve into the scaling law of parallel documents used for fine-tuning. Our exploration reveals that the performance of our LLM-based DocMT models shows considerable variation across diverse translation tasks. This variation ranges from surpassing gpt-4-turbo to achieving scores that are nearly zero. We focus exclusively on translations from English to German, Romanian, and Chinese, as our models demonstrate adequate proficiency in these specific translation tasks.
We showcase the results of the scaling law on English-German translation in Figure 4 and present the scaling law on English-Romanian and English-Chinese in Appendix F. It is observed that LLMs typically require only a minimal amount of training data for fine-tuning to reach high performance levels. However, various fine-tuning strategies exhibit distinct scaling behaviors. Specifically, our models employing LoRA require roughly $10\%$ of the training data (about $20K$ training examples) to match the performance of those fine-tuned on the entire training set. More impressively, our fully fine-tuned models need only about $1\%$ of the training data (approximately $2K$ training examples) to achieve performance nearly equal to that of the model fine-tuned on the full training set. Considering that the Llama2 variants are predominantly pre-trained on English text, these insights are particularly vital for low-resource languages.
En-De De-En
$d$BLEU COMET $d$BLEU COMET
Doc2Doc-mT5-1.2B 20.2 74.4 20.0 76.5
L-7B-LoRA 28.9 76.4 35.5 83.2
L-7B-fft 29.0 77.0 36.1 84.0
B-7B-LoRA 23.7 73.0 30.5 80.8
B-7B-fft 21.0 69.0 30.0 80.5
V-7B-LoRA 20.5 63.8 33.9 81.8
V-7B-fft 27.8 75.0 34.7 83.1
Table 4: $d$BLEU and COMET on WMT2023 test sets between English and German. Best results are highlighted in bold.
@@ 6.3 Scaling Law of Parallel Documents
6.4 Evaluation on Recent Test Sets
Given that LLMs are pre-trained on extensive text corpora, they are potentially susceptible to data leakage risks. To address this, we assess our models using additional, recent test sets from WMT2023 (Koehn et al., 2023). These tests, conducted between English and German, not only evaluate the out-of-domain generalization of our models but also help mitigate the risks associated with data leakage. The approaches in this work requires one-to-one sentence alignment in the parallel documents, but the parallel documents in the WMT2023 test sets are not structured in such a way. Hence, we utilize spaCy to segment the documents and discard any parallel documents where the source and target sides have a differing number of sentences. [^7] https://spacy.io/
Our findings are detailed in Table 4. It is evident that, while Doc2Doc-mT5 models surpass our LLM-based models as shown in Table 2, the LLM-based models significantly exceed the performance of the traditional Doc2Doc-mT5-1.2B on the recent out-of-domain test sets. These findings suggest that the LLM-based DocMT models are promising for translating out-of-domain text.
6.5 Zero-Shot Crosslingual Transfer
In this section, we investigate whether the translation capability acquired from translating between one language pair can be transferred to other language pairs. Specifically, we evaluate our English-German LLM-based DocMT models on the test sets from English to all the other languages.
We analyze the difference in COMET scores between our English-German LLM-based DocMT models and their backbones, as detailed in Table 5. Our findings reveal that models whose backbone is fine-tuned with instructions (Llama2-7B and Bloom-7B) exhibit positive transfer effects across all other language pairs. In contrast, models with an instruction-tuned backbone (Vicuna-7B) show limited improvement, benefiting only a few language pairs. These observations imply that during fine-tuning on parallel documents, LLMs are more likely to activate their inherent translation capabilities rather than developing new translation skills.
@@ 7 Conclusion
This study explores the adaptation of large language models (LLMs) for document-level machine translation (DocMT). The research involves extensive experimentation with two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. The results show that task-specific supervised fine-tuning on parallel documents significantly enhances the performance of moderately-sized LLM-based models ($7B$ parameters in size) in DocMT. These models achieve state-of-the-art results in specific translation tasks, even outperforming gpt-4-turbo sometimes. Furthermore, our study presents an in-depth analysis of the LLM-based DocMT models from three perspectives: translation error distribution, the scaling law of parallel documents, and the effects for zero-shot cross-lingual transfer. The findings of this research not only shed light on the strengths and limitations of LLM-based DocMT models but also provide a foundation for future advancements in the field of machine translation.
@@ Appendix A Statistics of Parallel Documents
Appendix B Prompt Types
@@ Appendix C GPT-4 Prompts
Appendix D Breakdown Results
@@
Appendix E Off-Target Translation
@@
Appendix F Scaling Law of Parallel Documents from English to Romanian and Chinese
In Section 6.3, we find that our LLM-based DocMT models are highly efficient in terms of the amount of training data. To confirm our findings in Section 6.3, we conduct additional experiments on the translation tasks from English to Romanian and Chinese. As shown in Figure 7, we can confirm the superiority of LLM-based DocMT models with regard to data efficiency.

### 15 Feb 2024 v2

@@ 5 Main Results
In our results presented in Table 2, we observe that gpt-4-turbo and gpt-3.5-turbo significantly outshine all other models in performance. Notably, the NLLB variants, which are trained on vast amount of parallel sentence pairs, also demonstrate superior performance among specialized machine translation (MT) models. In the context of DocMT, conventional DocMT models still outperform our LLM-based DocMT models for translations from English to other languages when evaluated using standard MT metrics. Conversely, for translations from other languages to English, our LLM-based DocMT models perform on par or better than conventional DocMT models in $\mu_{s\textsc{BLEU}}$ and $\mu_{d\textsc{BLEU}}$ metrics, while those conventional DocMT models maintain superior performance in $\mu_{\textsc{COMET}}$.
@@
As indicated in Table 2, our models incorporating LoRA typically outperform fully fine-tuned (FFT) LLMs. However, an exception is observed where V-7B-fft outperforms V-7B-LoRA in translating from other languages to English. This discrepancy is likely attributable to overfitting. In scenarios of extensive fine-tuning with a large corpus of parallel documents, the full fine-tuning of all parameters often leads to rapid overfitting on the training dataset. In contrast, the parameter-efficient fine-tuning approach, exemplified by LoRA, updates only a select number of parameters, effectively preventing the models from overfitting the training set. Furthermore, we observe that the L-7B and V-7B models exhibit comparable performance, suggesting that initializing with instruction-tuned models does not always enhance task-specific performance.
@@
We present the results for the translation tasks from other languages to English in Figure 2. Regarding the readability of the figures, we present only the results provided by our models using LoRA. Our LLM-based DocMT models exhibit superior performance, sometimes even surpassing gpt-4-turbo in certain translation tasks. However, they fail completely in others. A manual review of translation tasks where our LLM-based DocMT models fail reveals that the primary cause of failure is off-target translation. We provide an in-depth analysis of the off-target translation problem in Section 6. A complete breakdown of the results is in Appendix E.
002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8002,00020002{,}0002 , 0002,20022002{,}2002 , 2002,40024002{,}4002 , 400MistranslationOvertranslationUndertranslationAdditionOmissionGrammarCohesionCoherenceInconsistent styleMultiple terms in translation1,39513951{,}3951 , 3957177177177171,22412241{,}2241 , 2248408408408404844844844842342342342347097097097094544544544547257257257253463463463461,35613561{,}3561 , 3567177177177171,35813581{,}3581 , 3588638638638636166166166162662662662667277277277275655655655657717717717713393393393391,71217121{,}7121 , 7127157157157159779779779772452452452454174174174174864864864868048048048045925925925928248248248242732732732732,00220022{,}0022 , 0028368368368361,55115511{,}5511 , 551586586586586868868868868474474474474933933933933717717717717854854854854386386386386# of occurences
@@ 6 Analyses
In this section, we investigate the off-target problem and leverage gpt-4-turbo to analyze the translation errors. We also explore discourse phenomena, the training strategy, and the scaling law of parallel documents. Furthermore, we conduct additional evaluations on recent test sets from WMT2023 and examine crosslingual transfer.
@@
Off-Target Translation
In Figure 2, our LLM-based DocMT models excel in some translation tasks but struggle in others due to off-target translation issues. We investigate this problem using the fasttext library (Bojanowski et al., 2017) to identify translation languages and quantify off-target rates, which represent the proportion of translations that are off-target. Results are presented in Table 3, with off-target rates reaching up to $98.3\%$ in failing tasks. Notably, only B-7B-LoRA consistently maintains low off-target rates, likely due to Bloom-7B’s multilingual pre-training. These findings shed light on the main reason of translation failures in LLM-based DocMT models, offering insights for future research. Detailed off-target rates are provided in Appendix F.
Translation Errors
@@
To comprehensively understand the translation capabilities of our LLM-based DocMT models, we select specific error types from the Multidimensional Quality Metrics (MQM) framework (Burchardt, 2013). Kocmi and Federmann (2023) demonstrate gpt-4 is capable of identifying error spans and achieving state-of-the-art MT evaluation accuracy, so we leverage gpt-4-turbo to analyze the translation errors of the text translated by these models. We focus on four models due to resource constraints: L-7B-LoRA, L-7B-fft, Doc2Doc-mT5-1.2B, and GoogleTrans, assessing translations from English to German, Romanian, and Chinese. The error identification prompt is detailed in Appendix D, and we present the frequency of error types in Figure 3. Notably, most errors are limited to individual sentences. Despite similar scores in metrics such as $s$BLEU, $d$BLEU, and COMET among the models, our LLM-based DocMT models (L-7B-LoRA and L-7B-fft) exhibit fewer context-independent and context-dependent errors. This highlights a limitation in current evaluation metrics, suggesting they may not sufficiently assess document-level translations. It also indicates that fine-tuning LLMs for machine translation holds promise for enhancing DocMT performance.
Acc. er es sie
Doc2Doc-mT5-1.2B 77.0 68.7 89.0 73.5
MR-Doc2Sen-mT5 59.9 48.9 91.4 39.4
MR-Doc2Doc-mT5 78.2 67.5 91.1 76.1
DocFlat-mT5 78.0 68.9 90.1 75.1
IADA-mT5 79.1 70.0 89.8 77.6
L-7B-LoRA 83.1 77.2 96.6 75.4
L-7B-fft 81.1 70.2 96.9 76.2
B-7B-LoRA 75.5 56.2 95.1 75.1
B-7B-fft 68.3 50.8 95.5 58.5
V-7B-LoRA 84.9 78.4 96.2 80.1
V-7B-fft 84.4 76.3 96.4 80.5
Table 4: Accuracy (in %) on the English-German contrastive test set. Best results are highlighted in bold.
Discourse Phenomena
To evaluate our LLM-based DocMT model’s ability to leverage contextual information, we assessed it using the English-German contrastive test set by Müller et al. (2018). This evaluation tests the model’s accuracy in selecting the correct German pronoun (“er”, “es”, and “sie”) from multiple translation options. Results, shown in Table 4, reveal that models initialized with Llama2-7B and Vicuna-7B outperform Doc2Doc-mT5-1.2B, while Bloom-7B-initialized models perform worse, indicating that contextual understanding is mostly acquired during pre-training, as detailed by Scao et al. (2022) due to the lack of German text in Bloom pre-training.
$s$BLEU $d$BLEU COMET
Two-Stage
Nl-En 38.9 41.9 87.0
Ro-En 38.2 41.4 87.3
Ar-En 02.5 02.6 51.6
Zh-En 00.1 00.1 67.1
Three-Stage
Nl-En 39.1 42.1 87.0
Ro-En 38.4 41.6 87.3
Ar-En 02.3 02.4 52.4
Zh-En 00.3 00.3 67.4
Table 5: Comparison between two-stage and three-stage training strategies. The results of the two-stage strategy are given by L-7B-fft. For the three-stage training strategy, we fine-tune all the model parameters of Llama2-7B in all three stages.
1102030405060708090100404040406060606080808080Percentage (%) of training dataCOMET
@@
En-De De-En
$d$BLEU COMET $d$BLEU COMET
Doc2Doc-mT5-1.2B 20.2 74.4 20.0 76.5
MR-Doc2Sen-mT5 20.5 74.9 21.0 76.5
MR-Doc2Doc-mT5 21.2 75.6 21.5 76.5
DocFlat-mT5 20.9 75.1 21.8 76.5
IADA-mT5 21.2 75.4 22.0 76.5
L-7B-LoRA 28.9 76.4 35.5 83.2
L-7B-fft 29.0 77.0 36.1 84.0
B-7B-LoRA 23.7 73.0 30.5 80.8
B-7B-fft 21.0 69.0 30.0 80.5
V-7B-LoRA 20.5 63.8 33.9 81.8
V-7B-fft 27.8 75.0 34.7 83.1
Table 7: $d$BLEU and COMET on WMT2023 test sets. Best results are highlighted in bold.
Training Strategy
In this study, we follow the two-stage approach of Xu et al. (2023). Unlike traditional DocMT methods, which typically start with parallel sentence training, we explore the effectiveness of this conventional training strategy on LLM-based DocMT models. In this section, we introduce a three-stage training strategy, involving: (1) monolingual document fine-tuning, (2) parallel sentence fine-tuning, and (3) parallel document fine-tuning, for all parameters of the Llama2-7B. The results in Table 5 indicate that the three-stage training strategy is unnecessary for both high-performing languages (Dutch and Romanian) and low-performing languages (Arabic and Chinese) with LLM-based DocMT models.
Scaling Law of Parallel Documents
In this section, we explore the scaling law for fine-tuning parallel documents. We focus on English to German, Romanian, and Chinese translations due to our models’ proficiency. Results for English-German translation are presented in Figure 4, and for English-Romanian and English-Chinese in Appendix G. While LLMs typically excel with minimal training data, different fine-tuning strategies show distinct scaling behaviors. Our LoRA models match full training set performance with just $10\%$ of the data (around $20K$ examples), while fully fine-tuned models achieve near-equivalent performance with only about $1\%$ of the data (approximately $2K$ examples). These insights are crucial for low-resource languages, as recent LLMs are predominantly pre-trained on English text.
Evaluation on Recent Test Sets
Given their pre-training on extensive text corpora, LLMs may be susceptible to data leakage risks. We evaluate our models using recent test sets from WMT2023 (Koehn et al., 2023). These tests, conducted between English and German, not only evaluate the out-of-domain generalization of our models but also help mitigate the risks associated with data leakage. We use spaCy to segment documents and and discard any parallel documents where the source and target sides have a differing number of sentences. Our findings, presented in Table 7, reveal that while Doc2Doc-mT5 models outperform LLM-based models in Table 2, LLM-based models excel in translating out-of-domain text on the WMT2023 test sets. These findings highlight the ability of LLM-based DocMT to generalize well to out-of-domain translation tasks.
Zero-Shot Crosslingual Transfer
In this section, we explore the transferability of translation capabilities acquired from one language pair to others. We assess our English-German LLM-based DocMT models on English-to-other-language test sets, comparing their COMET scores to their base models in Table 6. Our results indicate that models with fine-tuned instructions (Llama2-7B and Bloom-7B) consistently exhibit positive transfer effects across all language pairs, while those with instruction-tuned backbones (Vicuna-7B) benefits only a few languages. These findings suggest that LLMs are more likely to activate their inherent translation abilities during fine-tuning rather than developing new ones.
@@ 7 Conclusion
This study investigates the adaptation of large language models (LLMs) for document-level machine translation (DocMT) through extensive experimentation with two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Results demonstrate that task-specific supervised fine-tuning on parallel documents significantly boosts the performance of moderately-sized LLM-based models (with $7B$ parameters) in DocMT, surpassing gpt-4-turbo in some cases. Our analysis offers insights into LLM-based DocMT models, providing a foundation for future advancements in the field of DocMT.
@@ Appendix A Statistics of Parallel Documents
Appendix B Optimization and Hyperparameters
Fine-tuning on Monolingual Documents
Fine-tuning on Parallel Documents
@@ Appendix B Optimization and Hyperparameters
Appendix C Prompt Types
@@ Appendix D GPT-4 Prompts
Appendix E Breakdown Results
@@
Appendix F Off-Target Translation
@@
Appendix G Scaling Law of Parallel Documents from English to Romanian and Chinese
In Section 6, we find that our LLM-based DocMT models are highly efficient in terms of the amount of training data. To confirm our findings in Section 6, we conduct additional experiments on the translation tasks from English to Romanian and Chinese. As shown in Figure 7, we can confirm the superiority of LLM-based DocMT models with regard to data efficiency.